OpenStudy (anonymous):

How do I solve this problem √(q-9)=7

OpenStudy (saifoo.khan):

Take square on both sides, (q-9) = 7^2

OpenStudy (anonymous):

start with \[q-9=7^2\] or \[q-9=49\] solve for q

OpenStudy (anonymous):

so 58?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks